x+y=12
x-y=4
slove using elimination

Answer by 303795(602)

If you add the first equation to the second one you will eliminate the y.
x + y + x - y = 12 + 4
2x = 16
so x = 8
Substitute x = 8 back into your original first equation
x + y = 12
so 8 + y = 12
y = 4
Check that x = 8, y = 4 works in the second equation
8 - 4 = 4
Correct